a projectile at any instant during its flight has velocity 5 m/s at 30° above the horizontal. how long after this instant will it be moving right angle to the given direction

Grade:11

1 Answers

Mohd Mujtaba
6 years ago
The time at which it be moving right angle to given vector is simply time of flight where u =initial velocity at given instant. So t=2usintheta /g=2×5×sin30/10=1/2=0.5sec. so answer is 《0.5sec.》 Thank ☺☺☺☺